OpenStudy (anonymous):

Find the solution, given the replacement set. 5x - 2y = 6 {(-2, -8), (0, -3), (1, -1), (2, 4), (4, 7) A.{(-2, -8), (2, 4), (4, 7)} B.{(-2, -8), (0, -3), (4, 7)} C.{(-2, -8), (0, -3), (1, -1)} D.{(1, -1), (2, 4), (4, 7)}

OpenStudy (amistre64):

so, which ordered pairs are solutions right?

OpenStudy (anonymous):

Yes I think so

OpenStudy (amistre64):

if so then we just plug them in to check :)

OpenStudy (amistre64):

does 1,-1 fit?

OpenStudy (anonymous):

I'd have to get out paper and pencil.

OpenStudy (amistre64):

paper and pencil are fine; but lets try this approach: 5(1) -2(-1) = 6

OpenStudy (amistre64):

does 5 + 2 = 6 ?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

then c and d are out

OpenStudy (amistre64):

a nad b differ by sat (0,-3) so if that fits its b, if not its a

OpenStudy (amistre64):

5(0)-2(-3) = 6 ? 0 + 6 = 6 ?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

a and b differ by say .... typing aint my forte :) then since b fits, lets use b :)

OpenStudy (amistre64):

since there is no "none of the above" option I assume it has to be one of those posted, so "b" seems the best candidate

OpenStudy (anonymous):

alright thanks!!